A company needs to pump coconut oil (den: 924 kg/m^3) from the import truck to its holding tank 20 meters above the ground. If the pump that is used is able to create 2 ATM of pressure, what will be the velocity of the oil as it enters the tank?
Answer #1

Solution :-

let the velocity of the oil is v

Using conservation of energy ,

Bernoulli's equation

P = p *g * h + 0.50 * p * v^2

2 * 1.01 *10^5 = 924 * 9.8 * 20 + 0.50 * 924 * v^2

solving for v

v = 6.72m/s

the velocity of oil in the tank is 6.72 m/s